# Follow-Bench: A Unified Motion Planning Benchmark for Socially-Aware Robot Person Following

### v2

@@ Abstract
Robot person following (RPF)—mobile robots that follow and assist a specific person—has emerging applications in personal assistance, security patrols, eldercare, and logistics. To be effective, such robots must follow the target while ensuring safety and comfort for both the target and surrounding people. In this work, we present the first end-to-end study of RPF, which (i) surveys representative scenarios, motion-planning methods, and evaluation metrics with a focus on safety and comfort; (ii) introduces Follow-Bench, a unified benchmark simulating diverse scenarios, including various target trajectory patterns, dynamic-crowd flows, and environmental layouts; and (iii) re-implements six popular RPF planners, ensuring that both safety and comfort are systematically considered. Moreover, we evaluate the two highest-performing planners from our benchmark on a differential-drive robot to provide insights into real-world deployment. Extensive simulation and real-world experiments provide quantitative insights into the safety-comfort trade-offs of existing planners, while revealing open challenges and future research directions. All code, data, and deployment scripts are publicly available at https://follow-bench.github.io/.
@@ I INTRODUCTION
Previous surveys on RPF[1, 3, 4] have provided broad overviews of its core components—such as perception, tracking, planning, control, and human-robot interaction—as well as factors[3] that influence human spatial preferences and expectations. Building on these foundations, this paper narrows the scope to the motion-planning dimension of RPF, with a particular emphasis on two critical and widely prioritized criteria: safety and comfort. Here, safety denotes continuous target tracking and reliable collision avoidance, while comfort primarily refers to maintaining appropriate interpersonal distance with respect to both the target and surrounding pedestrians. Guided by these requirements, we comprehensively review representative scenarios, planning methods, and evaluation metrics; propose a general RPF-driven motion-planning pipeline; and re-implement six popular planners to systematically analyze their performance.
@@
To enable systematic benchmarking, we introduce Follow-Bench, a unified benchmark for evaluating RPF planners under diverse conditions, including various target trajectory patterns, pedestrian-flow patterns, and environmental layouts. Unlike existing RPF benchmarks—which focus primarily on perception (e.g., identifying a specific target as in TPT-Bench[9] and Chen’s dataset[10]) or follow high-level language instructions in simplified environments (e.g., Uni-Navid[11] built on Habitat 3.0[12])—Follow-Bench targets the motion-planning capabilities required in realistic, socially interactive settings.
@@
Most existing RPF studies oversimplify human behavior by modeling pedestrians as randomly moving agents, neglecting realistic human interactions, and typically restrict evaluation to domestic or lab-scale environments with only a few pedestrians. Consequently, very few RPF approaches have been systematically evaluated in socially interactive scenarios, for evaluating the motion-planning capabilities for the purpose of accompanying the target person continuously and naturally while avoiding collision with others. For example, Repiso et al.[13, 14] evaluate their planners in a simulator that extends the social force model (SFM)[15] to account for interactions. However, their simulated scenarios are limited in scope and, critically, not released as an open-source unified benchmark.
@@
In this paper, we address these limitations by introducing Follow-Bench, which explicitly emphasizes two critical requirements for RPF: (i) continuously and naturally following a target person, and (ii) proactively avoiding obstacles and other pedestrians in socially interactive environments. Unlike independent navigation tasks[2, 16, 17, 18], which focus primarily on collision avoidance, these dual requirements underscore a fundamental trade-off between safety and comfort. Through extensive experiments with re-implemented popular planners in both simulation and real-world scenarios, we demonstrate how existing methods struggle to effectively balance this trade-off. In summary, the main contributions of this paper are:
@@
A unified benchmark, implemented in a lightweight simulator with realistic social interactions, for evaluating RPF motion-planning capabilities across diverse target trajectory patterns, pedestrian flows, and environmental layouts.
@@ II Literature Review
Comfort, in contrast, is more subjective and is typically reflected in the robot’s motion patterns and the relative distances between the robot and the target person and the surrounding pedestrians (i.e., proxemics [19]). Higher-level aspects of comfort are beyond the scope of this review, such as following and approaching behaviors [13], group following [20], or explicit human-robot interaction [21].
This review builds upon and complements prior surveys of RPF [1, 2, 3, 4] and related topics in social navigation [8, 22], but narrows the focus to representative scenarios (Sec. II-A) that challenge safety and comfort, and the evaluation metrics (Sec. II-B) used to assess performance with respect to these requirements. Furthermore, the reviewed methods (Sec. II-C) are analyzed with an emphasis on how they are designed to address these two key requirements.
@@ II-A Scenarios
We review representative RPF scenarios that pose challenges to both safety and comfort, encompassing both real-world and simulated environments. The difficulty of each scenario is influenced by multiple factors, including the layout complexity of the static environment, the intricacy of the target’s trajectory, and the degree of interference from dynamic objects (i.e., people). Accordingly, we categorize RPF scenarios along three dimensions: target trajectory pattern, dynamic-crowd flow and environmental scale and layout, as shown in Table I.
@@ II-A1 Target Trajectory Pattern
Some existing studies simplify the target trajectory pattern by modeling it as the obstacle-avoiding shortest path between predefined start and end points on a map. However, in RPF tasks, the target’s walking behavior and trajectory characteristics play a crucial role in determining the difficulty of following, introducing unique challenges for RPF control algorithms.
The complexity of a trajectory mainly arises from the target’s turning behaviors, which we categorize by difficulty into smooth turns and sharp turns. Smooth turns are typically associated with natural obstacle avoidance or casual walking patterns. These include standard L-shaped turns and square paths [7, 23, 24, 25, 26, 27], circular or figure-eight-shaped trajectories [7, 28], and consecutive S-shaped motions [29, 30, 27]. In contrast, sharp turns, which involve abrupt changes in direction, are particularly challenging for RPF systems. A common example is the U-turn [7, 28, 29, 30, 31, 32], though other forms such as obtuse-angle turns and triangular trajectories [7, 33] also pose significant difficulties. Leisiazar et al. [29] explicitly evaluated their RPF algorithm’s performance under varying target turning radii, highlighting the sensitivity of RPF systems to trajectory geometry.
The aforementioned turning behaviors typically involve simultaneous changes in both orientation and position during walking. However, in real-world scenarios, humans may adjust their orientation to a random direction. They would first reorient themselves—i.e., rotate in place—before proceeding in a different direction. This decoupling of orientation and translational movement has been explicitly considered in several RPF-related studies [29]. A notable distinction is made by Hu et al.[25] between step turns and spin turns, with the latter referring to in-place rotations that are particularly common during sharp directional changes. These behaviors introduce additional challenges for RPF systems, especially in target orientation estimation[34, 25] and following position adjustment to avoid obstructing the target’s intended path [7].
@@
| Dynamic-Crowd Flow | Pedestrian Number | One [38] |
@@
| Directionality and Pattern | Parallel Traffic [38, 39, 45, 46] | |
| Cross Traffic[39, 47, 45, 10] | ||
| Stuck [39, 44] | ||
| Perpendicular Traffic* | ||
| Join / Leave a Crowd* | ||
| Random Crowds* | ||
@@ II-A2 Environmental Scale & Layout
The scale and layout of an environment introduce distinct challenges to RPF. In small room-scale settings [31] or medium-sized single-floor spaces [35, 36], occlusions are frequent because visibility is limited and obstacles are dense; yet, once tracking is lost, re-finding the target is relatively straightforward thanks to the confined search area. Conversely, large-scale venues such as shopping malls [32] or open plazas exhibit fewer occlusions but make search and re-identification harder because the target can move towards a much broader region.
@@
Beyond scale, environmental layout adds layers of complexity. Narrow structures—hallways or corridors, for example—force the robot to keep a safe yet functional distance from both the target person and nearby pedestrians while navigating tight clearances [24, 31, 27, 38, 39]. When a wide space funnels into a doorway or passage, the follower often must switch its relative position to maintain line-of-sight and safety [13, 40]. Cluttered obstacles similarly demand a constantly shifting balance between obstacle avoidance and target pursuit [23, 40]. In addition, semantic cues embedded in the layout call for context-sensitive behavior: at intersections, the robot must anticipate cross-traffic and blind corners [37, 39]; in elevators, it should shrink its proxemic radius to respect the tighter personal space [8]; and in highly repetitive patterns—waiting-room bench grids, supermarket aisles, or office cubicles—it can exploit spatial regularity to plan more predictable and efficient paths. Finally, 3D topographical variations, such as ramps, stairs, and uneven terrains, introduce additional challenges related to stable human localization[41], locomotion and safe navigation[42]. Collectively, these factors underscore that successful RPF requires strategies tuned not only to environmental scale but also to the nuanced features of the surrounding layouts.
II-A3 Dynamic-Crowd Flow
@@
The complexity of RPF tasks increases in crowded scenes due to occlusion and safety risks caused by non-target pedestrians. Especially due to human-human interactions, pedestrian future motion is diverse and varied. In existing RPF simulated environments, some methods assume simplified crowd motion with constant velocity models [47, 33] without considering interactions, while some methods simulate the interactions via SFM or Extended SFM [45, 48, 44]. ORCA is another interaction simulation method, widely adopted in social navigation [8, 49]—shows promise for dense crowds, but its application to RPF remains unexplored.
@@
As for the pedestrian number settings, current evaluations primarily target sparse scenarios: Some studies focus on scenarios with only one other person[38] or medium-size groups (2-10 people) [43, 29, 10]. A few studies evaluate RPF performance in crowds with 20 people[44]. Although some scenarios involve many pedestrians, the people density of these scenarios is usually not provided. Besides, evaluations involving dense and dynamic crowds, like cafeterias and crossings, are limited.
@@
Furthermore, directionality and patterns of pedestrian flows, including crossing[39, 47, 45, 10], parallel passing[38, 39, 45, 46], and getting stuck in traffic[39, 44], also impact the robot’s following policy. For instance, the robot should leave a space for the target person to escape from getting stuck in crowds [44]. Apart from the above factors, additional target-crowd interactions should be considered, e,g., accompanying the target to cross a perpendicular traffic, and joining or leaving a crowd.
@@ II-B Evaluation Metrics
Safety and comfort are the two of the most frequently emphasized criteria in RPF research [8, 3]. Picture a target individual moving through an environment populated by static obstacles and other pedestrians. An ideal person-following robot should (i) avoid any collision and continuously observe the target, (ii) move smoothly, and (iii) remain in a socially acceptable zone relative to both the target and surrounding pedestrians. The core metrics discussed below are summarized in Table II.
In terms of safety, one common metric is the success rate of the following task, typically defined such that the task fails if more than one collision occurs [44]. Another key safety-related consideration is target visibility, especially important in scenarios with frequent occlusions. To this end, the target visibility ratio (TVR) is often used, measuring the proportion of time the robot successfully maintains tracking of the target [50, 51, 52]. When the target becomes occluded, the robot is expected to search and re-identify the target. In this context, the search success rate (SR) and the success weighted by path length (SPL) are commonly adopted [53]. SPL jointly considers both the success of person search and the efficiency of the path taken to complete the task.
Comfort is often assessed in terms of motion smoothness, typically evaluated through velocity, acceleration, and jerk [54, 55]. It also depends on the spatial relationship between the robot and the target. Following proxemic theory [19], previous works [33, 7] define a preferred region around the target, as illustrated in Fig. 1. Maintaining a position within the personal zone (0.45-1.2 meters) is generally considered desirable, and researchers often compute the duration the robot remains within this zone. In addition to distance preference, angle preference is also important (e.g., maintaining a side-following position). For instance, Repiso et.al [46] proposes a metric that measures the angular deviation from the preferred following angle.
Social awareness further requires maintaining appropriate distances not only from the target but also from surrounding pedestrians. Studies in social navigation [56, 57, 58] suggest maximizing the minimum distance to nearby individuals. For example, Arena [56] introduces a metric that quantifies the time spent within the private zone (less than 0.45 meters) of any pedestrian, with shorter durations indicating better social compliance. Path length is also frequently used as a proxy for energy efficiency [44, 56].
@@
| Category | Metric | Unit | Explanation | |||
@@
% | Runs with no collision | |||
@@
| Comfort | Path Length | [m] | Robot path length | |||
| Velocity (avg.) | [m/s] | Robot velocity | ||||
| Acceleration (avg.) | [m/s2] | Robot acceleration | ||||
| Movement Jerk | [m/s3] | Derivation of acceleration | ||||
| Time in personal zone | [s] |
| Time in private zone | [s] |

### v3

@@ Abstract
Robot person following (RPF)—mobile robots that follow and assist a specific person—has emerging applications in personal assistance, security patrols, eldercare, and logistics. To be effective, such robots must follow the target while ensuring safety and comfort for both the target and surrounding people. In this work, we present the first comprehensive study of RPF, which (i) surveys representative scenarios, motion-planning methods, and evaluation metrics with a focus on safety and comfort; (ii) introduces Follow-Bench, a unified benchmark simulating diverse scenarios, including various target trajectory patterns, crowd dynamics, and environmental layouts; and (iii) re-implements six representative RPF planners, ensuring that both safety and comfort are systematically considered. Moreover, we evaluate the two best-performing planners from our benchmark on a differential-drive robot to provide insights into real-world deployment of RPF planners. Extensive simulation and real-world experiments provide quantitative study of the safety-comfort trade-offs of existing planners, while revealing open challenges and future research directions. All code, data, and deployment scripts are publicly available at https://follow-bench.github.io/.
@@ I INTRODUCTION
Although prior surveys[1, 3, 4] have provided broad overviews of its core modules (e.g., perception, tracking and human-robot interaction) and social factors, a dedicated discussion of motion-planning principles that explicitly account for safety and comfort remains lacking. This paper addresses this gap by centering on these two fundamental criteria, which jointly define how robots can accompany humans in a socially acceptable and physically secure manner. Here, safety denotes continuous target tracking and reliable collision avoidance, while comfort primarily refers to maintaining appropriate interpersonal position with respect to both the target and surrounding pedestrians. Guided by these requirements, we comprehensively review representative scenarios, planning methods, and evaluation metrics, propose a general RPF-driven motion-planning pipeline, and re-implement six popular planners to systematically analyze their performance.
@@
To enable systematic performance evaluation, we introduce Follow-Bench, a unified benchmark for RPF planners under diverse conditions, including various target trajectory patterns, pedestrian-flow patterns, and environmental layouts. Unlike existing RPF benchmarks—which focus primarily on perception (e.g., identifying a specific target as [9, 10]) or follow high-level language instructions in simplified environments (e.g., [11] built on Habitat 3.0[12])—Follow-Bench targets the motion-planning capabilities required in realistic, socially interactive settings.
@@
Most existing RPF studies oversimplify human behavior by modeling pedestrians as randomly moving agents, neglecting realistic human interactions, and typically restrict evaluation to domestic or lab-scale environments with only a few pedestrians. Consequently, very few RPF approaches have been systematically evaluated in socially interactive scenarios, for evaluating the motion-planning capabilities for the purpose of accompanying the target person continuously and naturally while avoiding collision with others. For example, Repiso et al.[13, 14] evaluate their planners in a simulator that extends the social force model (SFM)[15] to account for interactions. Unfortunately, their simulated scenarios are limited in scope and, critically, are not released as an open-source benchmark.
@@
In this paper, we address these limitations by introducing Follow-Bench, which explicitly emphasizes two critical requirements for RPF: (i) continuously and naturally following a target person, and (ii) proactively avoiding obstacles and other pedestrians in socially interactive environments. Unlike independent navigation tasks[2, 16, 17, 18], which focus primarily on collision avoidance, these dual requirements underscore a fundamental trade-off between safety and comfort. Through extensive experiments with re-implemented representative RPF planners in both simulation and real-world scenarios, we demonstrate how existing methods struggle to effectively balance this trade-off. In summary, the main contributions of this paper are:
@@
A unified benchmark (Follow-Bench), implemented in a lightweight simulator with realistic social interactions, for evaluating RPF motion-planning capabilities across diverse target trajectory patterns, pedestrian flows, and environmental layouts.
@@ II Literature Review
Comfort, in contrast, is more subjective and is typically reflected in the robot’s motion patterns and the relative position between the robot and the target person and the surrounding pedestrians (i.e., proxemics [19]).
Guided by these two requirements, we comprehensively review representative scenarios (Sec. II-A) that challenge safety and comfort, and the evaluation metrics (Sec. II-B) used to assess performance with respect to these requirements. Furthermore, the reviewed methods (Sec. II-C) are analyzed with an emphasis on how they are designed to address these two key requirements.
@@ II-A Scenarios
RPF can occur in a wide range of real-world scenarios. Previous surveys [1, 3, 4] have typically categorized these scenarios by application domains or broad environmental contexts (e.g., indoor/outdoor, open/crowded spaces). However, such coarse classifications limit the ability to quantify scenario complexity, which is essential for benchmarking the capability of RPF planners. In practice, each scenario arises from a specific combination of factors—the target person’s walking behavior, the spatial layout of the environment, and the level of human interference surrounding the robot. Together, these factors determine how challenging it is for a robot to maintain both safety and comfort during following. To enable a more systematic description and comparison of existing studies, we categorize RPF scenarios along three key dimensions: target trajectory pattern, crowd dynamics, and environmental scale and layout, as summarized in Table I.
@@ II-A1 Target Trajectory Pattern
This dimension focuses on how the target person moves. In simple tests, the person follows a predefined shortest path on a map. In more realistic settings, however, their motion includes turns that increase the difficulty of following. Therefore, the primary source of trajectory complexity lies in the target’s turning behaviors, which we classify into smooth and sharp turns according to their difficulty.
Smooth turns are typically associated with natural obstacle avoidance or casual walking patterns. These include standard L-shaped turns and square paths [7, 20, 21, 22, 23, 24], circular or figure-eight-shaped trajectories [7, 25], and consecutive S-shaped motions [26, 27, 24]. In contrast, sharp turns, which involve abrupt changes in direction, are particularly challenging for RPF systems. A common example is the U-turn [7, 25, 26, 27, 28, 29], though other forms such as obtuse-angle turns and triangular trajectories [7, 30] also pose significant difficulties. Leisiazar et al. [26] explicitly evaluated their RPF algorithm’s performance under varying target turning radii, highlighting the sensitivity of RPF systems to trajectory geometry.
The aforementioned turning behaviors typically involve simultaneous changes in both orientation and position during walking. However, in some scenarios, humans may adjust their orientation to a random direction. They would first reorient themselves—i.e., rotate in place—before proceeding in a different direction. This decoupling of orientation and translational movement has been explicitly considered in several RPF-related studies [26]. A notable distinction is made by Hu et al.[22] between step turns and spin turns, with the latter referring to in-place rotations that are particularly common during sharp directional changes. These behaviors introduce additional challenges for RPF systems, especially in target orientation estimation[31, 22] and following position adjustment to avoid obstructing the target’s intended path [7].
@@
| Crowd Dynamics | Pedestrian Number | One [35] |
@@
| Directionality and Pattern | Parallel Traffic [35, 42, 43] | |
| Random Crowds [13, 14] | ||
| Perpendicular Traffic*[44, 36] | ||
| Circular Traffic*[44] | ||
| Join / Leave a Crowd*[18] | ||
@@ II-A2 Environmental Scale & Layout
This dimension concerns the environmental scale and layout in which RPF is performed. Small or medium-sized indoor spaces [28, 32, 33] often cause frequent occlusions due to dense obstacles and limited visibility, but re-finding the target is easier within a restricted area. Conversely, in large-scale venues such as shopping malls or open plazas [29], occlusions are fewer, yet the search and re-identification process becomes harder because the target can move towards a much broader region.
@@
In addition to scale, environmental layout is also a relevant factor in RPF complexity. Narrow structures—hallways or corridors, for example—force the robot to keep a safe yet functional distance from both the target person and nearby pedestrians while navigating tight clearances [21, 28, 24, 35, 36]. When a wide space funnels into a doorway or passage, the follower often must switch its relative position to maintain line-of-sight and safety [13, 37]. Cluttered obstacles similarly demand a constantly shifting balance between obstacle avoidance and target pursuit [20, 37]. In addition, semantic cues embedded in the layout call for context-sensitive behavior: at intersections, the robot must anticipate cross-traffic and blind corners [34, 36]; in elevators, it should shrink its proxemic radius to respect the tighter personal space [8]; and in highly repetitive patterns—waiting-room bench grids, supermarket aisles, or office cubicles—it can exploit spatial regularity to plan more predictable and efficient paths. Finally, 3D topographical variations, such as ramps, stairs, and uneven terrains, introduce additional challenges related to stable human localization[38], locomotion and safe navigation[39]. Collectively, these factors underscore that successful RPF requires strategies tuned not only to environmental scale but also to the nuanced features of the surrounding layouts.
II-A3 Crowd Dyanmics
@@
The complexity of RPF increases in crowded scenes due to occlusion and collision risks caused by non-target pedestrians. Especially due to human-human interactions, pedestrian future motion is diverse and varied. In existing RPF simulated environments, some methods assume simplified crowd motion with constant velocity models [46, 30] without considering interactions, while some methods simulate the interactions via SFM or Extended SFM [42, 45, 41]. ORCA is another interaction simulation method, widely adopted in social navigation [8, 47]—shows promise for dense crowds, but its application to RPF remains unexplored.
@@
As for the pedestrian number settings, current evaluations primarily target sparse scenarios. Some studies assume one other person[35] or a small group of 2-10 people [40, 26, 10]. A few studies evaluate RPF performance in crowds with 20 people[41]. Although some scenarios involve many pedestrians, the people density of these scenarios is usually not provided. Besides, evaluations involving dense and dynamic crowds, like cafeterias and crossings, are limited.
@@
Furthermore, directionality and patterns of pedestrian flows, including crossing[36, 46, 42, 10], parallel passing[35, 36, 42, 43], and getting stuck in traffic[36, 41], also impact the robot’s following policy. For instance, the robot should leave a space for the target person to escape from getting stuck in a crowd [41]. Apart from the above factors, additional target-crowd interactions should be considered, e,g., accompanying the target to cross a perpendicular traffic[36, 44, 10], and joining or leaving a crowd[18].
@@ II-B Evaluation Metrics
Safety and comfort are among the most frequently emphasized criteria in RPF research [8, 3]. Consider a scenario in which a target person moves through an environment populated with static obstacles and surrounding pedestrians. An ideal person-following robot should (i) avoid collisions and maintain continuous observation of the target, (ii) move with smooth and stable motion, and (iii) preserve a socially acceptable distance from both the target and nearby people. The core evaluation metrics corresponding to these criteria are summarized in Table II. Compared with the broadly reviewed metrics in prior surveys [3, 4], the metrics summarized here are streamlined to focus specifically on motion-planning-related aspects. Moreover, they are objectively quantifiable and categorized into two groups—safety-related and comfort-related metrics—to enable clearer capability analysis.
Regarding safety, a widely used metric is the avoidance success rate (ASR), which measures the percentage of trials completed without collisions [41]. Another key safety-related metric is the search success rate (SSR), defined as the percentage of times the robot successfully re-finds the target within a limited time after occlusions [48, 49, 50]. To further assess search efficiency, the success weighted by path length (SPL) metric is often adopted [48], jointly considering both SSR and the total path length taken during the search. Typically, ASR and SSR are combined into an overall success rate (SR), reflecting the overall completion performance of the RPF task.
Comfort is often assessed in terms of motion smoothness, typically evaluated through velocity, acceleration, and jerk [51, 52]. It also depends on the position between the robot and the target. Following proxemic theory [19], previous works [30, 7] define a preferred region around the target, as illustrated in Fig. 1. Maintaining a position within the personal zone (0.45-1.2 meters) is generally considered desirable, and researchers often compute the duration the robot remains within this zone. In addition to distance preference, angle preference is sometimes also important (e.g., maintaining a side-following position). For instance, Repiso et.al [43] proposes a metric that measures the angular deviation from the preferred following angle.
Moreover, the target visibility ratio (TVR) is another important comfort-related indicator, quantifying the proportion of time the robot successfully keeps the target within its field of view [53, 54, 49, 50]. The position between the robot and surrounding pedestrians also affects perceived comfort. Studies on social navigation [55, 56, 57] suggest maximizing the minimum distance to nearby individuals. For example, Arena [55] introduces a metric that measures the duration the robot spends within the private zone (less than 0.45 m) of any pedestrian, where shorter durations indicate better compliance with social norms. Finally, path length is frequently employed as a proxy for energy efficiency [41, 55].
@@
| Category | Metric | Unit | Explanation | ||||
@@
% | Runs with no collision | ||||
@@
| Comfort |
| Path Length | [m] | Robot path length | |||||
| Velocity (avg.) | [m/s] | Robot velocity | |||||
| Acceleration (avg.) | [m/s2] | Robot acceleration | |||||
| Movement Jerk | [m/s3] | Derivation of acceleration | |||||